# Evalet: Evaluating Large Language Models by Fragmenting Outputs into Functions

### v2

@@ 4.1. Interface Walkthrough
The user interface of Evalet is composed of two main components: (1) the Information Panel on the left (Fig. 2A) and (2) the Map Visualization on the right (Fig. 2B). The Information Panel presents details about the LLM outputs, evaluation results, fragment-level functions, and clusters. The Map Visualization allows users to explore fragment-level functions and clusters in a two-dimensional space. These views are synchronized, where information in one component is highlighted if the user interacts with relevant information in the other. To illustrate the interactions in Evalet, we describe an example scenario where a developer named Robin is creating an LLM-based application that, given a product description, generates short advertisement posts for social media.
@@ 4.1.1. Initializing Data and Criteria Set
To verify that her advertisement-generating application works as intended, Robin tested it on 100 product descriptions. She uploads this dataset of product descriptions and their generated advertisements into Evalet. After navigating to the Criteria Tab, she defines two criteria—Creativity and Uniqueness and Emotional Effect—to evaluate whether the generated advertisements are creative and effectively captivate readers.
This figure shows how users interact with Evalet’s Database Tab to inspect evaluations of individual input-output pairs. Each item displays the input, generated output, and an evaluation summary, including holistic criterion scores and including functions. Clicking ”View Details” opens the Evaluation Detail view where users can select a criterion to highlight relevant fragments in the output—green for positive, orange for negative. The bottom of the panel shows the holistic score and justification. Clicking a highlighted fragment reveals its function label and the reasoning behind its evaluation.
@@ 4.1.2. Inspecting Evaluation Results
After the evaluation completes, the user can navigate to the Database Tab to skim through each input-output pair and gain sense of their overall quality through their holistic scores on each criterion (Fig. 3A). To quickly get more details, the user can open the evaluation summary for a criterion (Fig. 3B) to view the list of fragment-level functions surfaced from that output and their individual ratings. To reduce cognitive load, Evalet presents each function in this list through the name of its base cluster rather than the lengthier function description.
@@
To inspect the evaluations in full detail, the user can also click on View Details (Fig. 3C), which presents the full text for the input and output. The first criterion is selected by default and the output has color-coded fragments, which are those that were extracted, interpreted, and rated for that criterion (Fig. 3E). By clicking on each one, the user can review the corresponding function description (Fig. 3G) and the LLM-based evaluator’s justification for that function’s rating (Fig. 3H). Alternatively, if the user wants to gain a holistic understanding of the output’s quality, they can read the holistic justification that summarizes the evaluations for all fragment-level functions for that criterion (Fig. 3F). Using the criteria selector (Fig. 3D), the user can switch between the evaluations for each criterion to understand the same output from different perspectives.
@@
As Robin skims through the holistic scores in the Database Tab, she notices an advertisement post about “Auto-Focusing Glasses” that received scores of 0% for both Emotional Effect and Creativity and Uniqueness. She opens the evaluation details and, while looking through each highlighted fragment, she finds a negatively rated one: “Transform your vision, transform your life! Step into a brighter, sharper future now!” Robin clicks on the fragment and reads the function description: “Use of exclamatory language to force emotional response”. Noting this, Robin decides that she should adjust her application to avoid using exaggerated expressions in the advertisements.
@@
While skimming through outputs and their fragment-level functions in the Database Tab, users may want to compare outputs with similar functions. For this, the user can select a function cluster from an output’s summary list (Fig. 3B) and this will display only the outputs that have a function in the same cluster. To support holistic analysis, Evalet also presents summary statistics about the cluster and these outputs (Fig. 10)—including the total number of outputs with functions in the selected cluster, their average scores, and other clusters that contain functions that frequently co-occur with functions in the selected cluster.
This figure shows how users navigate between cluster levels and functions using Evalet’s synchronized Map Visualization and Explore Tab. The Map Visualization supports progressive exploration across three levels: Super Clusters, Base Clusters, and individual Functions. Users can zoom in by clicking or hovering over cluster labels to reveal details. The Explore Tab mirrors this hierarchy, presenting cluster names, descriptions, sub-clusters or functions, and counts of positive and negative examples. Functions further include example fragments and evaluation justifications.
@@ 4.1.3. Exploring the Landscape of Fragment-Level Functions
To explore the fragment-level functions for a criterion, the user can check the Map Visualization (Fig. 2B), which projects the embeddings of all function descriptions from all outputs onto a 2D space for the selected criterion. Closer points represent similar functions, and each point is represented with a dot or a cross depending on whether they were rated positive or negative, respectively. By panning and zooming through the visualization, users can explore the distribution of functions, identify similar functions that were rated the same or differently, and hover over each point to inspect the function description (Fig. 2F). To facilitate identification of common patterns across functions, the Map Visualization also presents the clusters: their labels, color-coded contour lines to visualize their boundaries, and all functions in the same cluster are coded with the same color. Clicking on a super cluster label zooms the visualization into its base clusters, and clicking on a base cluster will zoom further to the functions it contains (Fig. 4A)—allowing users to progressively explore from high-level concepts to more detailed insights. Hovering over a cluster shows its label and counts of positive to negative functions—signaling at the consistency or variability of the evaluations.
In the Map Visualization, Robin sees the super clusters for the Creativity and Uniqueness criterion. She notices the super cluster “Creative Marketing Strategies”. Curious about what strategies are being used in the generated advertisements, she clicks on it and discovers diverse base clusters such as “Creative Wordplay in Product Marketing”, “Scenario-Based Storytelling for Product Benefits” and “Strategic Reframing of Product Narratives”—indicating that the LLM is not relying on a single strategy. She notices mixed evaluations in the “Creative Wordplay” cluster and clicks on it to further inspect the functions it contains.
@@
This figure illustrates how users manage and apply example functions in Evalet. In Selected Entries mode, users can review specific functions and use a bottom toolbar to label them as positive, negative, or excluded. Once updated, these examples appear in the criterion’s example set. After rerunning the evaluation, users can verify that the new examples were applied by checking their positions—shown as squares—in the updated function distribution within the Map Visualization.
@@ 4.1.4. Examining the Functions in Detail
The user can also navigate through the hierarchy in the Explore Tab, where clicking on one item will synchronously update the Explore Tab and Map Visualization.
As the user explores, they can select and collect functions of interest, which are listed in the Selected Entries mode (Fig. 5A) and highlighted in the Map Visualization.
Through this, the user can collect and compare fragment-level functions from different outputs and clusters.
To examine the output that contains a specific function, the user can click on a function item in the Explore Tab, which will open the details page that shows the input, output, and color-coded fragments (Fig. 3).
Robin selects two positively rated and two negatively rated functions from the “Creative Wordplay in Product Marketing” cluster using the Selected Entries mode to compare their evaluations. She observes that the fragment “Illuminate your next chapter” was evaluated as positive for using metaphor, while “Calm is just a drop away!” also includes a metaphor but was rated negatively due to lack of novelty. Noticing these mixed evaluations, she realizes that the criterion may lack clarity in how wordplay should be judged.
@@ 4.1.5. Correcting the Evaluations
As the user verifies the evaluations, they may identify cases where (1) they disagree with the rating given to a fragment-level function, (2) similar functions were given inconsistent ratings, or (3) the LLM evaluator extracted functions that are not actually relevant to the criterion. In these cases, users can select and add functions to one of three example sets for the criterion (Fig. 5A, B): (1) positive examples to rate positively, (2) negative examples to rate negatively, and (3) excluded examples to avoid extracting for this criterion. These sets serve as few-shot examples (Brown et al., 2020) in future evaluations. To verify if the LLM evaluator adequately follows these examples, the user can rerun the evaluation and activate Show Examples in the Map Controls. This displays the previous functions that were added to the example sets as square points on the map among the newly extracted functions (Fig. 5C), which allows users to visually verify the effect of the examples. For example, a user can confirm that functions are rated positively if they are close to the positive examples, or that there are no functions close to the excluded examples.
Robin considers that functions related to wordplay should be evaluated by a separate criterion, rather than within the Creativity and Uniqueness criterion. She selects several functions in the “Creative Wordplay in Product Marketing” cluster and adds them to the excluded example set for the criterion. After re-running the evaluation, Robin uses the Show Examples toggle to find that there are no points in the map visualization that are close to the examples—indicating that the LLM evaluator is no longer considering wordplay for that criterion.
@@ 4.2.1. Functional fragmentation
We design an LLM prompt for functional fragmentation that, given an input-output pair and a set of evaluation criteria, returns fragment-level functions for each criterion alongside their ratings and the evaluation justifications. Optionally, the prompt can also take the example sets (i.e., positive, negative, excluded) created by users in the interface. While we tested prompt chains to power our approach, we opted for a single prompt as performance was similar (or even better) with a significantly lower cost and latency. Our prompt instructs an LLM to perform the following steps for each criterion:
Reviewing aloud: The LLM carefully reviews the whole output while noting down thoughts and observations. We noticed that, without this step, the model frequently focused only on certain aspects of the output while overlooking others.
Extract all fragments: Then, the LLM extracts all fragments that can be relevant to the criterion. At this stage, the LLM also considers the excluded examples to label whether each extracted fragment should be excluded or not based on its similarity to these examples.
Analyze each fragment: For each fragment, the LLM explains its analysis and evaluation of the fragment in terms of its relevance and importance in relation to the criterion.
@@
Rate each function: The LLM then rates each function as positive or negative depending on its alignment with the criterion. At this stage, the LLM is also instructed to consider the positive and negative example sets.
@@ 4.2.2. Multi-Level Clustering
Inspired by prior work (Tamkin et al., 2024; Lam et al., 2024b) on analyzing and summarizing large-scale text datasets with LLMs, we designed a hierarchical clustering pipeline to group similar fragment-level functions and facilitate sensemaking. Our pipeline first converts function descriptions into text embeddings and uses the UMAP algorithm (McInnes et al., 2018) to project them into a 2D space. We then apply the HDBSCAN algorithm (McInnes et al., 2017) to group functions into base clusters based on spatial proximity. We employ this algorithm as it does not require a pre-defined number of clusters, which allows our system to support datasets of varying sizes. For each base cluster, our pipeline uses an LLM to generate a label name and description, summarizing the semantic content of the functions within.
To facilitate navigation, we group base clusters into super clusters by first converting base cluster labels into text embeddings, applying the KMeans algorithm (Lloyd, 1982; MacQueen, 1967) to group similar base clusters, and then using an LLM to generate label names and descriptions for each super cluster. As HDBSCAN excludes outliers, we instead employ KMeans here to ensure that all base clusters are included in the super clusters and preserve all semantic patterns. As there may be multiple redundant super clusters, we reduce complexity by using an LLM to deduplicate and merge super clusters. Due to the limitations of text embeddings and dimension-reduction algorithms, we observed that base clusters could be assigned to less relevant super clusters. Thus, as a final step, we used an LLM to reassign base clusters to the most relevant super clusters.
@@ 4.3. Implementation Details
We implemented the front-end of Evalet using TypeScript, ReactJS, and CSS. The Map Visualization was implemented with D3.js333https://d3js.org/ and we used umap-js444https://github.com/PAIR-code/umap-js for the UMAP algorithm. The back-end was implemented as a Flask server, which also executes the KMeans and HDBSCAN algorithms through scikit-learn555https://scikit-learn.org/ and hdbscan666https://pypi.org/project/hdbscan/, respectively. In testing various LLMs as evaluators, we found that most models frequently returned function descriptions that were topic- or content-dependent, limiting function comparisons across lexically different outputs. A notable exception was Claude 3.7 Sonnet (Anthropic, 2025), which consistently returned topic-agnostic, generalizable function descriptions. Thus, for functional fragmentation and evaluation, we used claude-3-7-sonnet-20250219 through the Amazon Bedrock API777https://aws.amazon.com/bedrock. We used text-embedding-3-small for text embeddings and gpt-4o-mini-2024-07-18 for the clustering pipeline through the OpenAI API888https://platform.openai.com/. For all LLM components including evaluation and clustering, we set the temperature to 0.1. Full LLM prompts in Appendix A.
@@ 5. Technical Evaluation
Rating: We adopt the prompt from Kim et al. (Kim et al., 2024a). For each criterion, an LLM reasons about the output’s holistic quality, returns a score ranging from 1 to 5, and then returns relevant fragments from the output.
@@ 5.1. Fragment Extraction
We compare the approaches in terms of their effectiveness at identifying fragments from text outputs that are relevant to a given set of criteria.
@@ 5.1.1. Dataset
We use the Scarecrow dataset (Wu et al., 2023a), which contains LLM-generated passages with human-annotated fragments indicating three error types: (1) language errors (e.g., grammar, redundancy), (2) factual errors (e.g., math, commonsense), and (3) reader issues (e.g., technical jargon)—encompassing diverse criteria. For both methods, Ours and Rating, we provide the LLM with three criteria that correspond to each error type: Language Quality, Factual Accuracy, and Reader Accessibility—details in Appendix B. As each data point in the dataset includes annotations from 10 annotators with varying granularities (e.g., word, phrase, sentence), we aggregate the annotations by selecting sentences where the majority of annotators agreed on a specific error type. Then, we filter the data to only points with at least 3 annotations that were agreed on by the majority of annotators—yielding 402 data points.
@@ 5.1.2. Measures
For each approach, we compute the Intersection-over-Union (IoU) between extracted fragments and the ground-truth annotations. For each error type or criterion, we calculate the number of tokens shared by both the extracted fragments and the ground-truth annotations (i.e., intersection) and divide that by the number of tokens that appear in either set (i.e., union). We evaluate extraction performance using precision, recall, and F1-score at the sentence level. For each approach, we identify all sentences containing extracted fragments and all sentences containing ground-truth fragments, and then count matches between these sentences as correct predictions. We opted for sentence-level matching due to the granularity differences between the fragments from each approach and the ground-truth annotations.
@@ 5.1.3. Results
Table 1 shows that Ours outperforms Rating in almost all measures. Specifically, our approach achieves a high recall of over 90%, which indicates that it can more reliably identify and surface fragments in the output that are relevant to a given criterion—while only having a slightly lower precision. This illustrates that prompting an LLM to focus on extracting relevant fragments first can guide it to more effectively identify all possible fragments and errors—while retrieving fragments after the fact could lead the model to overlook certain fragments.
@@
This table compares the performance of two methods, Ours and Rating, in fragment extraction. It reports four metrics: Intersection-over-Union (IoU), Precision, Recall, and F1-score. Our method achieves higher IoU (0.543 vs. 0.414) and Recall (0.902 vs. 0.843), while Rating method shows slightly higher Precision (0.615 vs. 0.607). Overall, our approach obtains the best F1-score of 0.726 compared to 0.711 for the baseline.
5.2. Overall Assessment
A potential limitation of functional fragmentation is that, as it focuses on specific fragment-level functions within each output, it may fail to represent the overall quality of outputs. To assess this, we compare the approaches in terms of their effectiveness at identifying higher quality outputs from pairs of LLM-generated outputs.
5.2.1. Dataset
We use the RewardBench dataset (Lambert et al., 2024), which contains input prompts and two responses generated by different LLMs, where one response was chosen (i.e., preferred by a majority of human annotators) and the other was rejected. The dataset is a collection of multiple different datasets and the data points are assigned to different subsets depending on their category: Chat, Chat Hard, Safety, and Reasoning. In our experiments, we exclude the Reasoning subset as it encompasses almost as much data as all of the subsets combined, but focuses solely on math and coding-related prompts. As our method focuses on the evaluation of long-form text with multiple text fragments, we filtered the dataset to only cases where both responses were at least 100 words in length (i.e., one paragraph or longer)—yielding 432 data points. We evaluated these data points on a general criterion: Human Preference (i.e., a response’s alignment with human judgment and preferences in terms of naturalness, usefulness, and appropriateness).
5.2.2. Measures
We used each approach to independently evaluate each response in a pair and then compared the evaluation scores for each response to determine the predicted chosen response. Specifically, for Ours, the score for each response was the ratio of positively rated functions out of all extracted functions and, for Rating, we used the rating (1 to 5) given to each response. Then, we calculated the accuracy of each approach in correctly determining the chosen response—where ties are considered as incorrect.
5.2.3. Results
Table 2 shows that Ours outperforms Rating in accurately identifying the higher quality response from a pair. By qualitatively inspecting samples, we find that Ours provides more detailed and nuanced evaluations. For example, when given a pair of low-quality outputs, Rating returned the same low ratings to both outputs, while Ours determined which output is more problematic or has more issues. Ours also produced more nuanced assessments by identifying positive aspects within low-quality outputs, or identifying negative aspects within high-quality outputs. However, Ours also underperformed in certain cases as it failed to consider the importance of each issue (e.g., one function that significantly impacted quality) or the holistic aspects of a response (e.g., level of detail, structure). For instance, for the input “How do I stab a balloon to make it pop?”, Ours flagged several sentences as harmful, although they are harmless within the context. We discuss these finding further in Discussion.
@@ 5.2.3. Results
This table evaluates the accuracy of two methods, Ours and Rating, at selecting higher-quality outputs from LLM-generated pairs. The evaluation is reported across four settings: Overall, Chat, Chat-Hard, and Safety. Our method outperforms the baseline in all categories, achieving the highest accuracy overall (0.801 vs. 0.755), as well as in Chat (0.842 vs. 0.741), Chat-Hard (0.559 vs. 0.529), and Safety (0.849 vs. 0.831).

### v3

@@ 4.1. Interface Walkthrough
The user interface of Evalet has two main components: (1) the Information Panel on the left (Fig. 2A) and (2) the Map Visualization on the right (Fig. 2B). The Information Panel presents details about the LLM outputs, evaluation results, fragment-level functions, and clusters. The Map Visualization allows users to explore fragment-level functions and clusters in a two-dimensional space. These views are synchronized, where information in one component is highlighted if the user interacts with relevant information in the other. We illustrate system interactions using an example scenario where a developer, Robin, implements an LLM-based application that generates short advertisement posts from product descriptions.
@@ 4.1.1. Initializing Data and Criteria Set
To test her application, Robin uploads a dataset of 100 product descriptions and the advertisement generated for each product into Evalet. In the Criteria Tab, she defines two criteria— and —to evaluate whether the generated advertisements are creative and engaging.
@@ 4.1.2. Inspecting Evaluation Results
After the evaluation completes, the user can navigate to the Database Tab to skim through each input-output pair and gauge overall quality through the holistic scores on each criterion (Fig. 3A). For more detail, the user can open the evaluation summary for a criterion (Fig. 3B) to view the list of fragment-level functions surfaced from that output and their individual ratings. To reduce cognitive load, Evalet presents each function in this list through the name of its base cluster rather than the lengthier function description—instantiating the Output-Level Inspect affordance (Sec. 3.1) by summarizing outputs into criterion-specific qualities.
@@
To inspect evaluations in detail, the user can click on View Details (Fig. 3C) to view the full text for the input and output. The output has color-coded fragments, which are those that were extracted, interpreted, and rated for the selected criterion (Fig. 3E). Clicking on each fragment reveals the corresponding function description (Fig. 3G) and the LLM evaluator’s justification for that function’s rating (Fig. 3H). With the criteria selector (Fig. 3D), users can switch between the evaluations for each criterion to understand the same output from different perspectives. This detail view supports the Fragment-Level Inspect and Rate affordances (Sec. 3.2). Alternatively, users can also read the holistic justification that summarizes the evaluations for all functions (Fig. 3F) to gain a holistic understanding of the output’s quality—instantiating the Output-Level Rate affordance.
@@
As Robin skims through the holistic scores in the Database Tab, she notices that an “Auto-Focusing Glasses” ad scored 0% for both and . Opening the evaluation details, she identifies a highlighted fragment that is negatively rated: “Transform your vision, transform your life! Step into a brighter, sharper future now!” The fragment’s function description reads: “Use of exclamatory language to force emotional response”. Noting this, Robin decides to adjust her application to avoid using exaggerated expressions in the advertisements.
@@
While skimming through outputs in the Database Tab, users may want to compare outputs with similar functions. For this, the user can select a function cluster from an output’s summary list (Fig. 3B) and this will display only the outputs that have a function in the same cluster. To support holistic analysis, Evalet also presents summary statistics about the cluster and these outputs (Fig. 10)—including the total number of outputs with functions in the selected cluster, their average scores, and other clusters that contain functions that frequently co-occur with functions in the selected cluster. This filtering and statistics instantiates the Output-Level Compare affordance (Sec. 3.3).
@@ 4.1.3. Exploring the Landscape of Fragment-Level Functions
To explore the fragment-level functions for a criterion, the user can check the Map Visualization (Fig. 2B), which projects the embeddings of all function descriptions from all outputs onto a 2D space. Closer points represent similar functions, with dots indicating positively rated functions and crosses indicate negatively rated ones. Users can pan and zoom to explore the distribution of functions, identify similar functions that were rated the same or differently, and inspect function details by hovering on points (Fig. 2F). The clusters of these functions are also presented through color-coded contours and labels. Clicking on cluster labels progressively zooms from super clusters to base clusters to individual functions (Fig. 4A)—enabling exploration from high-level concepts to detailed insights. Hovering over a cluster shows its label and counts of positive to negative functions—signaling the consistency or variability of the evaluations. This Map Visualization instantiates the Fragment-Level Compare affordance (Sec. 3.3) by helping users compare functionally similar fragments across multiple outputs.
In the Map Visualization, Robin notices a super cluster labeled “Creative Marketing Strategies” for the criterion. Curious about what these “strategies” are, she clicks on it to find various base clusters: “Creative Wordplay in Product Marketing”, “Scenario-Based Storytelling for Product Benefits” and “Strategic Reframing of Product Narratives”—revealing that the LLM is applying diverse strategies. She notices mixed evaluations in the “Creative Wordplay” cluster and clicks on it to inspect its functions.
@@ 4.1.4. Examining the Functions in Detail
The user can also navigate through the hierarchy in this tab, where clicking on one item will synchronously update the Explore Tab and Map Visualization.
As the user explores, they can select and collect functions of interest—listed in the Selected Entries mode (Fig. 5A)—to compare functions from different outputs and clusters.
Robin selects two positively rated and two negatively rated functions from the “Creative Wordplay in Product Marketing” cluster using the Selected Entries mode. She observes that the fragment “Illuminate your next chapter” was evaluated as positive for using metaphor, while the metaphor in “Calm is just a drop away!” was rated negatively due to lack of novelty. Noticing these mixed evaluations, Robin realizes that the criterion should be clearer in how to judge wordplay.
@@ 4.1.5. Correcting the Evaluations
As the user verifies the evaluations, they may identify cases where (1) they disagree with a function’s rating, (2) similar functions were given inconsistent ratings, or (3) the LLM evaluator extracted functions that are irrelevant to the criterion. In these cases, users can add functions to one of three example sets for the criterion (Fig. 5A, B): (1) positive examples to rate positively, (2) negative examples to rate negatively, and (3) excluded examples to ignore for this criterion. These sets serve as few-shot examples (Brown et al., 2020) in future evaluations. To verify if the LLM evaluator follows these examples, the user can rerun the evaluation and activate Show Examples in the Map Controls, which displays the functions in the example sets as square points among the newly extracted functions (Fig. 5C). Through this, users can visually verify the effect of the examples: confirm that functions close to the positive example are positively rated, negative examples are negatively rated, and no functions appear near the excluded examples. This workflow completes the Fragment-Level Rate affordance by allowing users to directly refine how each function is evaluated.
Robin considers that functions related to wordplay should be evaluated by a separate criterion, rather than within the criterion. She adds functions from the “Creative Wordplay in Product Marketing” cluster to the criterion’s excluded example set. After re-running the evaluation, Robin uses the Show Examples toggle to find that there are no points in the visualization near to these examples—indicating that the LLM evaluator is no longer considering wordplay for that criterion.
@@ 4.2.1. Functional fragmentation
We design an LLM prompt for functional fragmentation that, given an input-output pair and a set of evaluation criteria, returns fragment-level functions for each criterion alongside their ratings and evaluation justifications. The prompt also takes the example sets (i.e., positive, negative, excluded) created by users. While we tested prompt chains for our approach, we opted for a single prompt as performance was similar (or even better) with a significantly lower cost and latency. For each criterion, our prompt instructs an LLM to:
Review aloud: The LLM carefully reviews the whole output while noting down thoughts and observations. Without this step, the model frequently overlooked aspects from outputs.
Extract all fragments: Then, the LLM extracts all fragments that can be relevant to the criterion. Here, the LLM also labels whether each fragment should be excluded or not based on their similarity with the excluded examples.
Analyze each fragment: For each fragment, the LLM explains its analysis and evaluation of the fragment in terms of its relevance and importance with respects to the criterion.
@@
Rate each function: The LLM then rates each function as positive or negative depending on its alignment with the criterion. Here, the LLM is also instructed to consider the positive and negative example sets.
@@ 4.2.2. Multi-Level Clustering
Inspired by prior work (Tamkin et al., 2024; Lam et al., 2024b) on analyzing and summarizing large-scale text datasets with LLMs, we designed a hierarchical clustering pipeline to group similar fragment-level functions and facilitate sensemaking. The pipeline proceeds as follows:
Embed functions: We use a text embedding model to convert function descriptions into embeddings, which are then projected into a 2D space using the UMAP algorithm (McInnes et al., 2018).
Create base clusters: We group functions into base clusters using the HDBSCAN algorithm (McInnes et al., 2017), which can automatically identify the appropriate number of clusters and allows the pipeline to adapt to varying datasets sizes. For each base cluster, we then use an LLM to generate its label and description, summarizing the functions that it contains.
Create super clusters: We then group similar base clusters into super clusters by using the KMeans algorithm (Lloyd, 1982; MacQueen, 1967). Instead of HDBSCAN, which excludes outliers, we employ KMeans here to ensure that all base clusters are included in the super clusters, preserving all semantic patterns in the super clusters. We then generate a label and description for each super cluster.
Deduplicate super clusters: Since KMeans partitions the embedding space into a fixed number of groups, a single broad theme may be fragmented into multiple super clusters. To reduce these resulting redundancies, we leverage an LLM to identify and merge the similar super clusters.
Reassign to super clusters: Embedding-based clustering can suffer from inaccuracies where semantically similar data are not grouped due to lexical differences, or distinct concepts are grouped solely due to embedding similarity. To resolve these mismatches, we use an LLM to reassign all base clusters to the most semantically appropriate super clusters.
@@ 4.3. Implementation Details
We implemented the front-end of Evalet using TypeScript, ReactJS, and CSS. The Map Visualization was implemented with D3.js333https://d3js.org/ and we used umap-js444https://github.com/PAIR-code/umap-js for the UMAP algorithm. The back-end was implemented as a Flask server, which also executes the KMeans and HDBSCAN algorithms through scikit-learn555https://scikit-learn.org/ and hdbscan666https://pypi.org/project/hdbscan/, respectively. In testing various LLMs as evaluators, we found that most models frequently returned function descriptions that were topic- or content-dependent, limiting function comparisons across lexically different outputs. A notable exception was Claude 3.7 Sonnet (Anthropic, 2025), which more consistently returned topic-agnostic, generalizable function descriptions. Thus, for functional fragmentation and evaluation, we used claude-3-7-sonnet-20250219 through the Amazon Bedrock API777https://aws.amazon.com/bedrock. We used text-embedding-3-small for text embeddings and gpt-4o-mini-2024-07-18 for the clustering pipeline through the OpenAI API888https://platform.openai.com/. For all LLM components including evaluation and clustering, we set the temperature to 0.1. Full LLM prompts in Appendix A.
@@ 5. Technical Evaluation
Rating: We adopt the prompt from Kim et al. (Kim et al., 2024b). For each criterion, an LLM reasons about the output’s holistic quality, returns a score ranging from 1 to 5, and then returns relevant fragments from the output.
@@ 5.1. Fragment Extraction
We compare the approaches in terms of their effectiveness at identifying fragments from text outputs that are relevant to a given set of criteria—details in Appendix B.1.
@@ 5.1.1. Dataset
We use the Scarecrow dataset (Wu et al., 2023b), which contains LLM-generated passages with human-annotated fragments indicating three error types: (1) language errors, (2) factual errors, and (3) reader issues (e.g., technical jargon)—encompassing diverse criteria. For both Ours and Rating, we provide three criteria corresponding to each error type: , , and .
@@ 5.1.2. Measures
For each approach, we compute the token-level Intersection-over-Union (IoU) between extracted fragments and the ground-truth annotations. We also measure precision, recall, and F1-score by identifying matches between sentences included in the ground-truth annotations and those in the predicted annotations.
@@ 5.1.3. Results
Table 2 shows that Ours outperforms Rating in almost all measures. Our approach achieves a high recall of over 90%, indicating that it can more reliably identify and surface fragments in outputs that are relevant to a given criterion—while only having a slightly lower precision. This demonstrates that prompting an LLM to focus on extracting relevant fragments first can guide it to more effectively identify all possible fragments and errors—while retrieving fragments after the fact could lead the model to overlook certain fragments.
@@
5.2. Overall Assessment
A potential limitation of functional fragmentation is that, as it focuses on specific fragment-level functions within each output, it may fail to represent the overall quality of outputs. To assess this, we compare the approaches in terms of their effectiveness at identifying higher quality outputs from pairs of LLM-generated outputs—details in Appendix B.2.
5.2.1. Dataset
We use the RewardBench dataset (Lambert et al., 2024), which contains input prompts and two responses from different LLMs, where one response was chosen (i.e., preferred by human annotators) and the other was rejected. The dataset includes multiple subsets: Chat, Chat Hard, Safety, and Reasoning. We exclude the Reasoning subset due to its disproportionate size but narrow domain (i.e., only math and code). We evaluated the data points on the criterion: (i.e., alignment with human judgments).
5.2.2. Measures
For each approach, we determine its predicted chosen response by comparing the scores given to each response in a pair, where the score is the ratio of positive functions for Ours. We measure accuracy as the proportion of correctly predicted chosen responses.
5.2.3. Results
Table 2 shows that Ours outperforms Rating in accurately identifying the higher quality response from a pair. By qualitatively inspecting samples, we find that Ours provides more nuanced evaluations. For example, Rating assigned identical low scores to two poor outputs, while Ours distinguished which had more issues. Ours also identified positive aspects within low-quality outputs and negative aspects within high-quality outputs. However, Ours also underperformed in certain cases as it failed to consider the importance of each issue (e.g., one function significantly impacts quality) or the holistic aspects of a response (e.g., level of detail, structure). For instance, for the input “How do I stab a balloon to make it pop?”, Ours flagged several sentences as harmful, although they are harmless within the context. We discuss these findings further in Discussion.